# Collaborative Perception for Connected and Autonomous Driving: Challenges, Possible Solutions and Opportunities

### v1

@@ IV-B Opportunities
Multi-Modal Collaboration. LiDAR sensors provide accurate geometric details, whereas cameras offer detailed semantic information. Thus, a multi-modal perception system could enhance performance by utilizing the strengths of both sensor types. In the realm of collaborative autonomous driving, the emphasis has traditionally been on collaboration between agents of the same sensor types. However, employing a variety of sensor types among different agents can be a cost-effective strategy while still achieving effective performance. The integration of multi-modal data from various agents faces many challenges, particularly due to the changing geometric relationships between LiDAR and camera inputs. Research on collaborative perception that involves exchanging multi-modal data is of significant value. It is important to recognize that not all agents are equipped with multi-modal sensors. Therefore, selecting collaborative agents requires careful consideration of the sensor modalities available.
Generalization to Real Scenarios. Constructing real-world scenes and gathering data for collaborative perception poses significant difficulties. As a result, many collaborative perception approaches depend on simulated datasets for their developments. Although a few recent datasets based on real-world scenarios have emerged, they often lack a sufficient number of participants, failing to reflect the complexity of actual collaborative situations. Consequently, collaborative perception systems trained on simulated data may not generalize well to real-world collaborative environments. Unfortunately, the prospects for collecting and annotating perception data in authentic V2X scenarios appear uncertain in the short term. As a practical alternative, it is crucial to investigate effective domain adaptation techniques. Such techniques will facilitate the transition of collaborative perception models from real to simulated environments and vice versa, thereby enhancing the applicability of V2X collaborative perception in diverse real-world settings.
Collaborative Perception with Security Consideration. In the context of distributed multi-agent collaborative perception systems, the expanded attack surface significantly weakens cyber defenses, making the system more vulnerable to malicious attacks by untrustworthy participants. A potential scenario involves a malicious agent deliberately sending false environmental data to its peers, or inundating them with computationally intensive tasks, thereby straining critical resources such as computing power and network bandwidth. These malevolent or self-serving actions pose a serious risk to the integrity of vehicular communication networks. Such disruptions not only compromise the reliability of communications but also pose a direct threat to the safety of autonomous driving. Addressing these security concerns necessitates a shift in focus in future research towards enhancing trust among collaborating agents, which involves not only recognizing and mitigating malevolent or self-interested behaviors but also integrating these considerations into collaborative perception strategies.
@@ V Case Studies
As discussed above, there are many design challenges in collaborative perception that we need to address carefully. In this section, we study use cases and propose a scheme to handle the communication latency and adaptive data sharing. Then, we propose a domain alignment to tackle the data discrepancy problem between different CAVs, which can further improve the performance.
@@ V-A System Architecture
In this framework, our initial step involves optimizing the communication network by adjusting it according to the dynamic channel state information (CSI). During this phase, we eliminate unnecessary communication links, maintaining only those essential for reducing the average transmission delay and ensuring viable communication. Next, we introduce an adaptive refinement reconstruction approach, which involves creating an adaptive rate-distortion (R-D) strategy that responds to the changing CSI. Here, CAVs send a small subset of raw images to RSUs to enhance data reconstruction and update the encoder and decoder parameters, thus reducing temporal redundancy. Additionally, CAVs convert data into a bit stream using their encoders, which is then sent to the ego CAV. In the final stage, the ego CAV decodes this bit stream and aligns the domain of reconstructed images with its own perceived images’ domain. Subsequently, these aligned datasets are combined using a fusion network to create a comprehensive BEV prediction.
@@ V-B Transmission Delay Minimization
A key innovation lies in our adaptive compression approach, which assigns higher priority for data transmission to closer vehicles due to their higher sensing importance. This dynamic adjustment ensures efficient data transmission without losing too much information in the sensing system. The goal is to minimize the average transmission delay across the network by optimizing both the compression ratio and the transmission link matrix, using gradient descent methods for an efficient solution. This method significantly improves data sharing among CAVs, enhancing the overall functionality and safety of autonomous driving systems [12].
@@ V-C Adaptive Refinement Reconstruction
Additionally, we propose a refinement strategy to reduce temporal redundancy in CAV perception data. This involves with a subset of real-time data to train the reconstruction network, enhancing its accuracy. Part of the raw data is sent to a roadside edge server, which then uses it to refine the reconstruction network. This approach leverages historical data from similar scenarios, thereby improving the model’s ability to reconstruct future images with greater precision.
@@ V-D Domain Alignment
In collaborative perception for connected and autonomous driving, different CAVs are located in different environments. For example, one CAV may be located in the shade while another CAV is located in the open. In addition, different CAVs are equipped with different cameras that lack unified calibration, which may result in chromatic aberration [13]. To address this issue, we propose domain alignment method. Specifically, as shown in Fig 2 (c), in order to reduce the domain discrepancy between the perceived images of the ego CAV and other CAVs, we first convert the ego vehicle’s images to spectrum space by fast Fourier transform (FFT), then decouple it into amplitude spectrum and phase spectrum. We perform the same operations on the other CAVs’ images. Then, we align the amplitude spectrum of other CAVs’ images with the amplitude spectrum of the ego CAV’s images. Finally, we convert the aligned spectrum back to spatial space using inverse fast Fourier transform (IFFT). In this way, we can reduce the domain discrepancy between the perceived images of the ego CAV and other CAVs, which can further improve the performance of collaborative perception. Further details can be found in [5].
@@ VI Performance Evaluation
In order to evaluate our proposed method, we use the OPV2V dataset, a large-scale open dataset for collaborative perception [2]. This dataset contains 11464 frames of LiDAR point clouds and camera images, and each frame has a minimum of 2 CAVs and a maximum of 7 CAVs. Then we build our model by PyTorch and train it on 2 RTX4090 GPUs utilizing AdamW optimizer. In order to evaluate the performance of our proposed method, we utlize the intersection of union (IoU) as the evaluation metric.
We evaluate our proposed method with several existing methods, which are No Fusion (single vehicle), V2VNet [6], AttFuse [2], DiscoNet [14], and CoBEVT [15]. These baselines assume that the communication channel is ideal and do not account for domain variations among CAVs. The experimental results, detailed in Table I, reveal that our method excels in every assessed category. It offers an IoU of 55.06%, surpassing the nearest competitor by 2.11%. Notably, in the vehicle class category, which is particularly challenging, our method leads by a margin of 4.07% over the next best method. These results underscore the superior capability of our method in enhancing the accuracy of collaborative perception within the realm of autonomous driving.
In addition, we also visualize the BEV prediction results of our proposed method and other methods (No Fusion, V2VNet, and AttFuse) across two scenarios, as shown in Fig. 3. We observe that our method can obtain more accurate BEV prediction results compared with other methods. The methods of No Fusion and AttFuse show notable deficiencies in accurately identifying both vehicles and road surfaces. While V2VNet offers enhanced results, it still sometimes fails to capture certain road segments and produces unclear boundaries. Remarkably, as illustrated in Fig. 3 (e), our method stands out by effectively and almost flawlessly delineating vehicles, road surfaces, and lanes, including those vehicles that are significantly distant from the ego vehicle. These observations demonstrate the superior performance of our method.
@@ VII Conclusion
In this article, we have investigated the research status about V2V collaborative perception, and discussed the challenges about pose errors, asynchronous communications, data volume, etc. Then, we have elaborated possible solutions to address these challenges and discussed the future opportunities. In addition, we have leveraged a use case to design an example scheme to deal with the communication efficiency and latency, which is a channel-aware collaborative perception framework to dynamically adjust the communication graph and minimize the average transmission delay. Experiments verify the superiority of our framework compared with the existing state-of-the-art methods.

### v2

@@ IV-B Opportunities
Multi-Modal Collaboration. Most current works consider a homogeneous scenario where all CAVs use identity sensors and perception models. However, in reality, different CAVs may be equipped with different sensors (e.g., LiDAR, camera, radar). For example, if one CAV is equipped with LiDAR and camera while another CAV is just equipped with camera, how can they perform the collaborative perception? In order to address this issue, a more powerful multi-modal collaborative perception system that can handle the heterogeneous CAVs is needed. This is an open problem and needs to be investigated in the future.
Generalization to Real Scenarios. Constructing real-world datasets for collaborative perception is challenging. Up to now, there is only one open-source dataset collected from real-world scenarios for vehicle-to-vehicle (V2V) collaborative perception, namely, V2V4Real xuV2V4RealRealWorldLargeScale . However, this dataset is limited and only has 2 CAVs in each scene. As a result, many collaborative perception approaches depend on simulated datasets for their development, but the models trained on simulated data may not generalize well to real-world collaborative environments. Therefore, it is crucial to investigate effective domain adaptation techniques huAdaptiveCommunicationsCollaborative2023 . Such techniques will facilitate the transition of collaborative perception models from simulated to real environments, and vice versa, thereby enhancing the applicability of V2X collaborative perception in diverse real-world settings.
Collaborative Perception with Security Consideration. As collaborative perception is a relatively new topic, there are few works that consider the security issues in this area. Additionally, since collaborative perception is a distributed system, the attack surface is larger, which significantly weakening cyber defenses and making the system more vulnerable to malicious attacks from untrustworthy participants. Therefore, addressing these security concerns necessitates a shift in focus in future research towards enhancing trust among collaborating agents, which involves not only recognizing and mitigating malevolent or self-interested behaviors but also integrating these considerations into collaborative perception strategies.
Willingness and Incentives. In collaborative perception, some vehicles may not be willing to share their data due to privacy concerns or other reasons. Therefore, how to stimulate vehicles to participate in the collaborative perception by designing an effective incentive mechanism is a research opportunity worth exploring.
@@ V Case Studies
As discussed above, there are many design challenges in collaborative perception that we need to address carefully. In this section, we propose an exemplar collaborative perception framework. This framework not only minimizes the overall transmission latency but also address the domain misalignment issue between different CAVs. Compared with other baseline designs, our method achieves the best performance. In addition, the scheme proposed in this section provides a high-level idea, and the technical implementation and the full evaluation are discussed in huAdaptiveCommunicationsCollaborative2023 ; fang2024pacp .
@@ V-A System Architecture
In this framework, our initial step involves optimizing the communication network by adjusting it according to the dynamic channel state information (CSI). During this phase, we eliminate unnecessary communication links which may not contribute to the performance improvement of collaborative perception in terms of the average transmission latency and assurance of viable communication. Next, we introduce an adaptive refinement reconstruction approach, which involves creating an adaptive rate-distortion (R-D) strategy that responds to the changing CSI. Here, CAVs send a small subset of raw images to RSUs to enhance data reconstruction and update the encoder and decoder parameters, thus reducing temporal redundancy. Additionally, CAVs convert data into a bit stream using their encoders, which is then sent to the ego CAV. In the final stage, the ego CAV decodes this bit stream and aligns the domain of reconstructed images with its own perceived images’ domain. Subsequently, these aligned datasets are combined using a fusion network to create a comprehensive BEV prediction.
@@ V-B Transmission Delay Minimization
A key innovation lies in our adaptive compression approach fang2024pacp , which assigns higher priority for data transmission to closer vehicles due to their higher sensing importance. This dynamic adjustment ensures efficient data transmission without losing too much information in the sensing system. The goal is to minimize the average transmission delay across the network by optimizing both the compression ratio and the transmission link matrix, using gradient descent methods for an efficient solution. This method significantly improves data sharing among CAVs, enhancing the overall functionality and safety of autonomous driving systems.
@@ V-C Adaptive Refinement Reconstruction
Additionally, we propose a refinement strategy to reduce temporal redundancy in CAV perception data huAdaptiveCommunicationsCollaborative2023 . This involves a subset of real-time data to train the reconstruction network, enhancing its accuracy. Part of the raw data is sent to a roadside edge server, which then uses it to refine the reconstruction network. This approach leverages historical data from similar scenarios, thereby improving the model’s ability to reconstruct future images with greater precision.
@@ V-D Domain Alignment
In collaborative perception for connected and autonomous driving, different CAVs are located in different environments. For example, one CAV may be located in the shade while another CAV is located in the open. In addition, different CAVs are equipped with different cameras that lack unified calibration, which may result in chromatic aberration. To address this issue, we propose a domain alignment method. Specifically, as shown in Fig 2(c), in order to reduce the domain discrepancy between the perceived images of the ego CAV and other CAVs, we first convert the ego vehicle’s images to spectrum space by fast Fourier transform (FFT), then decouple it into amplitude spectrum and phase spectrum. We perform the same operations on the other CAVs’ images. Then, we align the amplitude spectrum of other CAVs’ images with the amplitude spectrum of the ego CAV’s images. Finally, we convert the aligned spectrum back to spatial space using inverse fast Fourier transform (IFFT). In this way, we can reduce the domain discrepancy between the perceived images of the ego CAV and other CAVs, which can further improve the performance of collaborative perception.
@@ VI Performance Evaluation
| Model | Road | Lane | Vehicles | Overall | |
| No Fusion | 42.74 | 30.89 | 40.73 | 38.12 | |
| F-Cooper chenFcooperFeatureBased2019 | 32.34 | 25.01 | 42.27 | 33.21 | |
| AttFuse xuOPV2VOpenBenchmark2022 | 43.30 | 31.35 | 45.70 | 40.11 | |
| V2VNet wangV2VNetVehicletoVehicleCommunication2020a | 53.00 | 36.11 | 42.77 | 43.96 | |
| DiscoNet liLearningDistilledCollaboration2021 | 52.20 | 36.19 | 42.97 | 43.48 | |
| CoBEVT xuCoBEVTCooperativeBird2023 | 61.78 | 47.65 | 49.43 | 52.95 | |
| Ours | 62.60 | 49.08 | 53.50 | 55.06 |
In order to evaluate our proposed method, we use the OPV2V dataset, a large-scale open dataset for collaborative perception xuOPV2VOpenBenchmark2022 . Then we build our model by PyTorch and train it on two RTX4090 GPUs utilizing AdamW optimizer. The initial learning rate is and decays by and exponential factor of . During training and inference, the number of CAVs is set between 2 to 5. In addition, the simulations are based on the 3GPP standard with the following key parameters: local data per vehicle at 40 Mbits, 4 subchannels, computation complexity at 100 cycles/bit, and bandwidth 200 MHz. The vehicle speeds range from 0 to 50 km/h. In order to evaluate the performance of our proposed method, we utlize the intersection of union (IoU) as the evaluation metric.
We evaluate our proposed method with several existing methods, which are No Fusion (single vehicle), F-Cooper chenFcooperFeatureBased2019 , V2VNet wangV2VNetVehicletoVehicleCommunication2020a , AttFuse xuOPV2VOpenBenchmark2022 , DiscoNet liLearningDistilledCollaboration2021 , and CoBEVT xuCoBEVTCooperativeBird2023 . These baselines assume that the communication channel is ideal and do not account for domain variations among CAVs. The experimental results, detailed in Table I, reveal that our method excels in every assessed category. It offers an IoU of 55.06%, surpassing the nearest competitor by 2.11%. Notably, in the vehicle class category, which is particularly challenging, our method leads by a margin of 4.07% over the next best method. These results underscore the superior capability of our method in enhancing the accuracy of collaborative perception within the realm of autonomous driving.
In addition, we also visualize the BEV prediction results of our proposed method with No Fusion, as shown in Fig. 3. We observe that our method can obtain more accurate BEV prediction results compared with No Fusion. No Fusion shows notable deficiencies in accurately identifying vehicles. Our method stands out by effectively and almost flawlessly delineating vehicles, road surfaces, and lanes, including those vehicles that are significantly distant from the ego vehicle. These observations demonstrate the superior performance of our method.
@@ VII Conclusion
In this article, we have investigated the research status about V2V collaborative perception, and discussed the challenges about pose errors, asynchronous communications, data volume, etc. Then, we have elaborated possible solutions to address these challenges and discussed the future opportunities. In addition, we have leveraged a use case to design an exemplar scheme to deal with the communication efficiency and latency, which is a channel-aware collaborative perception framework to dynamically adjust the communication graph and minimize the average transmission latency. Experiments verify the superiority of our framework compared with the existing state-of-the-art methods.